OpenStudy (anonymous):

Solve for x. x/7 = -3

OpenStudy (anonymous):

x/7 = -3 -3*7 = -21 x = -21

OpenStudy (anonymous):

haha how so?

OpenStudy (anonymous):

its funny

OpenStudy (anonymous):

:P get over it. haha